# Selective Preference Optimization via Token-Level Reward Function Estimation

### v1

@@ 3.3 Application: Weak-to-Strong Generalization
We propose to leverage weak oracle models that underperform the policy models to select the key tokens from strong in-distribution data. Our intuition is that weak supervisors (oracle models) only identify which tokens are most effective in enhancing the alignment performance, rather than directly providing supervision signals, which normally requires stronger capabilities than the student model.
@@
During alignment, as the policy model becomes stronger, continual full optimization on the original data distribution can lead to over-optimization on the reward function Gao et al. (2023); Rafailov et al. (2024a), which can seriously degrade policy model performance. Online preference optimization Xiong et al. (2024); Xie et al. (2024) alleviates over-optimization with online annotations of new in-distribution data, but can be costly for strong policy models.
Weak data supervision focuses on scenarios when only weak out-of-distribution data is available for strong policy models. We propose to leverage the SePO framework to select key tokens from the weak dataset, and only the selected tokens are utilized to supervise the strong policy model. Instead of full optimization on the training data, we expect selective optimization on the key tokens to prevent over-optimization on the out-of-distribution data, while still leveraging effective supervision signals to further improve the policy model.
Arena-Hard AlpacaEval 2.0 MT-Bench
Policy Model Methods Win Rate LC Win Rate Win Rate Assistant QA Math Coding Avg.
Pythia-SFT-2.8B Base 2.34% 3.8% 4.12% 4.7 5.05 1.27 1.8 2.8
+DPO 5.71% 5.72% 6.1% 4.7 5.57 1.55 2.13 3.26
+IPO 5.6% 4.8% 4.96% 5.16 5.1 1.1 2.41 3.28
+RRHF 4.37% 4.33% 4.47% 4.5 4.82 1.6 2.25 2.93
+SimPO 5.2% 5.8% 6.0% 4.35 6.06 1.52 2.0 3.3
+SePO (Ours) 6.3% 7.1% 7.32% 5.3 5.3 2.38 2.5 3.45
Pythia-SFT-6.9B Base 4.23% 5.0% 5.17% 6.0 6.18 1.48 1.75 3.68
+DPO 10.2% 12.78% 13.27% 6.33 7.26 2.87 2.85 4.7
+IPO 8.1% 11.78% 12.6% 6.7 6.53 2.5 2.43 4.34
+RRHF 7.47% 11.42% 13.2% 6.46 6.28 2.5 2.4 4.31
+SimPO 8.0% 11.8% 12.72% 6.65 5.89 3.13 2.83 4.51
+SePO (Ours) 10.94% 14.27% 13.6% 6.7 7.37 2.93 2.85 4.79
TinyLLaMA-Chat Base 1.6% 1.26% 1.43% 4.01 3.48 1.45 2.35 3.28
+DPO 2.2% 1.95% 2.18% 4.02 3.52 1.5 2.4 3.31
+IPO 2.3% 1.85% 2.06% 4.03 3.72 1.45 2.4 3.38
+RRHF 3.1% 3.02% 2.12% 4.12 3.7 1.3 2.5 3.4
+SimPO 2.61% 1.3% 3.7% 4.02 4.07 1.35 2.25 3.28
+SePO (Ours) 4.1% 3.55% 3.37% 4.52 3.93 1.35 2.25 3.62
LLaMA2-Chat-7B Base 4.6% 5.4% 5.0% 6.48 7.87 1.45 1.95 4.48
+DPO 8.5% 7.8% 6.71% 6.33 6.26 2.85 2.85 5.43
+IPO 8.12% 8.78% 9.4% 6.66 6.98 2.55 2.3 5.75
+RRHF 9.4% 13.35% 14.41% 6.15 6.16 3.05 2.8 5.35
+SimPO 9.59% 13.58% 15.4% 6.52 6.38 3.5 2.8 5.63
+SePO (Ours) 10.3% 14.4% 14.91% 7.15 7.53 2.85 3.1 6.27
LLaMA2-Chat-13B Base 12.0% 8.4% 7.7% 5.82 6.36 3.15 3.3 5.7
+DPO 13.48% 13.72% 13.37% 6.92 6.81 4.1 3.77 5.84
+IPO 13.95% 14.27% 14.4% 7.02 6.63 4.3 3.5 5.76
+RRHF 13.84% 15.94% 16.36% 5.85 6.38 3.7 3.25 5.73
+SimPO 14.7% 16.4% 17.02% 6.73 6.4 3.2 3.0 5.7
+SePO (Ours) 15.5% 17.53% 18.41% 7.2 6.9 3.55 3.95 6.36
Table 1: Performance of SePO and other baseline methods on three benchmark datasets. For SePO, the oracle models are based on TinyLLaMA-1.1B and Pythia-1B, trained on the full UltraFeedback dataset. The modeled reward function is then used to select the top-30% tokens of chosen and rejected responses.
@@ 4.1 Experimental Settings
Models and Training Data
We comprehensively evaluate SePO on two representative model families: LLaMA (Touvron et al., 2023) and Pythia Biderman et al. (2023). To approximate the optimal token-level reward function, we first obtain the base model by training on the UltraChat-200K dataset Ding et al. (2023) in an SFT manner. For the LLaMA model family, we train the base model on the pre-trained TinyLLaMA-1.1B Zhang et al. (2024) model. For Pythia, we separately train on Pythia-(70M, 160M, 410M, 1B, 1.4B) to facilitate research on the effect of oracle models with different sizes. For each base model, we obtain the oracle models by further fine-tuning with DPO on the UltraFeedback dataset Cui et al. (2023) with 61,135 preference pairs. We examine the selective preference training process on target policy models with various capabilities. For Pythia, we perform SFT on Pythia-(2.8B, 6.9B) with UltraChat-200K and use them as policy models, which we refer as Pythia-SFT. For LLaMA, we test on TinyLLaMA-Chat-1.1B, LLaMA2-Chat-7B, and LLaMA2-Chat-13B.
Baseline Methods
We compare the performance of SePO with state-of-the-art offline preference optimization methods to indicate its effectiveness. DPO Rafailov et al. (2024c) explicitly models the relation between the reward functions and optimal policy, which enables reward model-free preference alignment. IPO Azar et al. (2024) provides a variant of DPO that prevents the requirement that pairwise preferences can substitute reward values. We introduce another two competitive reference model-free alignment methods: RRHF Yuan et al. (2024) directly optimize the probability of the target response pairs with a pairwise ranking loss, with an SFT-based regularization on the chosen responses. SimPO Meng et al. (2024) optimizes the length-regularized probability of the response pairs with a margin.
Evaluation Benchmarks
We quantify the contribution of each method by evaluating on three widely utilized instruction-following benchmarks: AlpacaEval 2.0 Dubois et al. (2024), MT-Bench Zheng et al. (2024b), and Arena-Hard Li et al. (2024). AlpacaEval 2.0 consists of 805 queries to evaluate the models’ versatile conversational abilities. We report the win rates and length-controlled win rates (LC win rate) of evaluated models against the responses of GPT-4-turbo. The LC win rates are designed to reduce influences of model verbosity. MT-Bench covers eight categories with 80 queries. We report the average scores ranging from 0 to 10. Due to the widely reported poor separability of MT-Bench reported by previous works Meng et al. (2024); Li et al. (2024), we further display fine-grained scores of model capability, which we reorganize the original 8 categories as follows: Assistant (Writing, Roleplay, Extraction), QA (Reasoning, STEM, Humanities), Math, and Coding. Scores for Assistant and QA are averages of the original categories. Arena-Hard extends MT-Bench with 500 high-quality queries, where we report win rates against GPT-4-0314 model outputs. All judgments are performed by the latest GPT-4o model.
@@ 4.2 Overall Performance
Performances of SePO and other baseline methods on three benchmark datasets are presented in Table 1. According to the results, SePO significantly improves performance over the base policy models. For example, on Arena-Hard, SePO achieves a 6.71% improvement in win rates on Pythia-6.9B. SePO also outperforms other strong baseline preference optimization methods. On Arena-Hard, SePO achieves the best win rates on all policy models, surpassing state-of-the-art methods such as SimPO and DPO. As SePO models are only optimized on 30% of the tokens trained on other methods, these results directly strengthen the effectiveness of selective training strategies applied in preference optimization.
On AlpacaEval 2.0, SePO continues to achieve superior performances over baseline methods on both win rates and LC win rates, further proving its effectiveness on different benchmarks. Notably, SePO outperforms all other methods on length-controlled win rates, including SimPO and RRHF, which are specifically designed for length-normalized reward formulation. These results show that selective training strategies also enhance policy models in avoiding over-length responses. We believe the reason is that during token selection, the token-level reward function can assign the end-of-sentence tokens with low-ranking scores, which can be discarded during optimization if the response ends inappropriately (e.g. over-length or repetition). In contrast, though SimPO and RRHF design length-normalized rewards, the end-of-sentence tokens are included for all responses, still fitting the response lengths for all training samples.
On MT-Bench, SePO outperforms all other methods on average scores. Due to the widely discussed poor separability of overall scores for MT-Bench, we look into category-based evaluations that provide fine-grained assessments. As shown, SePO achieves the best performances on 70% of comparisons on Assistant and QA, indicating its significant improvement on subjective tasks that require high-level intention understanding and writing skills. However, SePO outperforms baseline methods in math and coding on only 40% of the comparisons, underperforming baseline methods such as IPO and SimPO on several policy models. A possible reason is that objective tasks such as math and coding require coherent logic along the token-level MDP for response generation Xie et al. (2024); Chen et al. (2024b); Lai et al. (2024), while SePO is only optimized on selected tokens, which brings discontinuity in learning the logic during training. Baseline methods that optimize all tokens enable policy models to learn the full chain of reasoning and show advantages in objective scenarios.
Figure 4: SePO performances with different combinations of Top-Kselection ratios for chosen and rejected responses. The performances are quantified by the LC win rates on the AlpacaEval 2.0 benchmark. All results are obtained with a TinyLLaMA oracle model trained on the full UltraFeedback dataset.
@@ 4.3 Impact of Data Scale
According to the results, increasing selection rates from 0.1 for chosen/rejected responses rapidly improves policy model performance, but the momentum decreases as the ratio continues to grow. For example, the LC win rate of LLaMA2-Chat-7B improves from 8.37% to 14.8% as the ratios for chosen/rejected responses rise from 0.1 to 0.5 progressively, then stabilizes around 14.7% with higher selection rates. These observations prove our hypothesis that not all tokens are equally effective for LLM alignment. Training only on key tokens effectively improves alignment performance, while other tokens provide limited supervision information. From the figures, we conclude that training on Top-30% tokens for TinyLLaMA or Top-50% tokens for LLaMA2-Chat-(7B, 13B) already provides comparable performance to alignment on all tokens.
Increasing selection rates for chosen responses generally outperforms increasing ratios for rejected responses. For example, with a fixed rejected selection rate, the TinyLLaMA-Chat performance smoothly improves as the chosen ratios grow. However, improving rejected ratios from 0.1 to 0.9 leads to decreased model performance in 4 out of 5 fixed chosen selection rates. Similar results can be observed in the other two policy models. These results indicate that compared to increasing probabilities for irrelevant tokens in chosen responses, suppressing probabilities for high-reward tokens in rejected responses can significantly affect the model performance.
Figure 5: SePO performances supervised by oracle models trained with different proportions of the target dataset. The performances are quantified by the LC win rates on the AlpacaEval 2.0 benchmark. For each proportion, we do random sampling three times and report the average performance.
@@
In Theorem 2, we proved that training an oracle model on a randomly selected moderate-scale dataset is a pessimistic estimation of the target reward function. In this section, we empirically investigate the influence of the training data scale for oracle models on SePO performance. Specifically, we randomly sample different proportions of data from UltraFeedback as training data for the TinyLLaMA-based oracle model. The results are shown in Figure 5.
@@
Figure 6: Evaluation results on weak-to-strong generalization. (a) LC win rates of policy models on AlpacaEval 2.0, which are trained with oracle models of various sizes; (b)(c) Token-level score distributions for 5,000 chosen/rejected responses, with the rewards obtained from different oracle models.
@@ 4.4 Weak-to-Strong Generalization
In Table 1, we presented the performance of TinyLLaMA and Pythia-1B oracle models on guiding stronger policy models (e.g. Pythia-SFT-6.9B, LLaMA2-Chat-13B). The competitive results of SePO prove the viability of weak oracle modeling. To provide a clear landscape, we further train oracle models with Pythia-(70M, 160M, 410M, 1B, 1.4B) on the same target data and compare their performances on the Pythia-SFT-(2.8B, 6.9B) policy models. The results are shown in Figure 6 (a).
According to the results, oracle models with weak capabilities can provide effective supervision to strong policy models. For example, training with the Pythia-410M oracle model achieves 6.58% on Pythia-SFT-2.8B and 13% on Pythia-SFT-6.9B policy models, with up to 16.8$\times$ more parameters than the oracle model. These results outperform full optimization on the target dataset with baseline methods such as DPO and SimPO. In addition, the performance of target policy models continually improves as the oracle model size increases. For example, on Pythia-SFT-6.9B policy model, the 1.4B oracle model outperforms the 410M orcale model by 1.84% and the 70M model by 9.15%. These results show that oracle models with stronger capabilities can better model the token-level reward function and accurately select key tokens.
To provide an intuitive view, we present the token-level score distributions of different oracle models in 6 (b)(c). For the chosen scores distribution, strong oracle models such as Pythia-(1B,1.4B) show higher densities in extreme (large and small) reward values, which facilitates separating key tokens from the other tokens. In contrast, small oracle models tend to fit a Gaussian distribution, where most tokens have similar scores. These results show that weak oracle models struggle to distinguish key tokens within the dataset. We observe similar results on rejected scores distribution, which further proves the capability of the oracle models as crucial in modeling the token-level reward function.
@@
Table 2: Performance of SePO and other baseline methods on generalizing the weak HH-RLHF dataset to the strong LLaMA2-Chat-13B policy model.
Weak Data Supervision.
We evaluate the weak data supervision performance of SePO by training on HH-RLHF Bai et al. (2022), an early-released preference dataset with relatively lower quality Yang et al. (2024a) on responses. We perform SePO with a TinyLLaMA-based oracle model and 30% token selection rates, and comparisons with baseline methods are shown in Table 2. According to the results, SePO is the only method that improves the strong LLaMA2-Chat-13B policy model with data from HH-RLHF, outperforming base performance by 1.63% on Arena-Hard and 0.41% on AlpacaEval 2.0. With full optimization, baseline methods such as DPO and SimPO continuously degrade model performance due to over-optimization on weak supervision data. These results prove SePO effective in leveraging useful supervision signals from weak data while avoiding fitting harmful patterns. These results point to SePO as a highly efficient method for continually improving strong model performance with large-scale out-of-distribution data.
5 Related Work
5.1 Response-Level Preference Optimization
Aligning large language models (LLMs) with human preferences has been a key research focus, traditionally achieved through Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020). While effective, RLHF methods often face challenges like training instability and inefficiency, motivating the development of more direct alignment strategies. Recent approaches have emerged to address these issues without relying on complex reward modeling. Rafailov et al. (2024c) introduce Direct Preference Optimization (DPO), a method that simplifies RLHF by extracting the optimal policy in closed form, enhancing stability and reducing computational demands. Yuan et al. (2023) propose Reinforcement Ranking from Human Feedback (RRHF), which aligns model outputs through a ranking loss, streamlining the training process and minimizing the need for extensive hyperparameter tuning. Meng et al. (2024) further simplify alignment with Simple Preference Optimization (SimPO), using an implicit reward based on average log probability to improve both computational efficiency and performance. Ethayarajh et al. (2024) present KTO, which integrates human biases from prospect theory into the alignment process, achieving superior results by maximizing the utility of model outputs. Lu et al. (2024) propose SamPO, a technique that addresses verbosity in DPO by mitigating length-based biases in reward estimation, leading to significant performance gains. Finally, Azar et al. (2024) introduces a general framework that improves upon RLHF and DPO by eliminating common approximations, offering both theoretical insights and empirical benefits.
5.2 Token-Level Preference Optimization
Response-level direct alignment methods can be suboptimal due to the sequential, auto-regressive nature of token generation in LLMs. This has led to the development of token-level alignment methods using Markov Decision Processes (MDP) Rafailov et al. (2024b); Zhong et al. (2024); Zeng et al. (2024); Yoon et al. (2024); Chen et al. (2024c); Chan et al. (2024). Rafailov et al. (2024b) extend DPO to a token-level MDP, enabling better alignment with LLMs’ auto-regressive structure and improving credit assignment. Zhong et al. (2024) introduce Reinforced Token Optimization (RTO), combining token-level rewards with DPO and PPO for more efficient policy learning. Zeng et al. (2024) propose Token-level Direct Preference Optimization (TDPO), which optimizes policy at the token level using forward KL divergence, enhancing alignment and diversity. Yoon et al. (2024) present Token-Level Continuous Reward (TLCR), which assigns continuous rewards to tokens based on contextual confidence, improving alignment. Chen et al. (2024c) develop ALLO, a method that reduces overfitting by focusing on optimizing the most relevant neurons for alignment. Chan et al. (2024) use attention weights to redistribute rewards across tokens, stabilizing training and improving performance. Additionally, Chen et al. (2024a) propose DICE (DPO ImpliCit rEwards), a self-alignment method that uses implicit DPO rewards to further refine alignment without external feedback. Further applications of these token-level methods have been explored in the context of mathematical reasoning Lai et al. (2024); Chen et al. (2024b); Xie et al. (2024); Setlur et al. (2024), where they demonstrate potential in improving the precision and coherence of LLMs on complex reasoning tasks.
5.3 Weak-to-Strong Generalization
Weak-to-strong generalization, where weaker models supervise stronger ones, has become a significant topic in the ongoing development of LLMs Burns et al. (2023); Lang et al. (2024); Yang et al. (2024b); Charikar et al. (2024). This approach addresses the challenge of aligning increasingly powerful models with human values, particularly as models surpass human-level capabilities. Burns et al. (2023) show that strong models fine-tuned on labels from weaker supervisors can outperform their weak teachers, though naive fine-tuning has limitations and may not scale well with superhuman models. Lang et al. (2024) introduce a framework explaining how strong models can correct weak models’ errors and generalize beyond their knowledge, enhancing learning from incomplete or incorrect labels. Yang et al. (2024b) discuss the risk of "weak-to-strong deception" where strong models exploit weak supervisors to appear aligned while misbehaving in unmonitored areas, stressing the need for more robust alignment strategies. Charikar et al. (2024) quantify the performance gains of strong models over weaker ones, introducing misfit error as a key metric for optimizing this process. Additional studies have applied weak-to-strong generalization in tasks like high-quality token selection Lin et al. (2024), weak-to-strong search Zhou et al. (2024), and alignment techniques such as Aligner Ji et al. (2024) and Weak-to-Strong Extrapolation Zheng et al. (2024a).
6 Discussion
Conclusion.

### v2

@@ 3.3 SePO for Weak-to-Strong Generalization
We propose to leverage weak oracle models to select key tokens from in-distribution data. Our intuition is that weak supervisors (oracle models) only identify which tokens are most effective in alignment, rather than directly providing supervisions, which normally requires stronger capabilities than student models.
@@
When only weak out-of-distribution data is available, we propose to leverage SePO to select key tokens from the weak dataset, and only the selected tokens are utilized to supervise the strong policy model. Oracle models trained on high-quality data are used to directly perform key token selection on OOD data. We expect selective optimization to prevent over-optimization on the out-of-distribution data, while still leveraging effective supervision signals to further improve the policy model.
Arena-Hard AlpacaEval 2.0 MT-Bench Arena-Hard AlpacaEval 2.0 MT-Bench
Methods Policy Model GPU Hours Win Rate LC Win Rate Win Rate GPT-4o Policy Model GPU Hours Win Rate LC Win Rate Win Rate GPT-4o
Base Pythia-2.8B 2.34% 3.8% 4.12% 2.8 LLaMA2-Chat-7B 4.6% 5.4% 5.0% 4.48
+DPO 79.45 5.71% 5.72% 6.1% 3.16 182.95 8.5% 7.8% 6.71% 5.43
+IPO 75.9 5.6% 4.8% 4.96% 3.12 170.3 8.12% 8.78% 9.4% 5.64
+RRHF 52.19 4.37% 4.33% 4.47% 2.93 123.67 9.4% 13.35% 14.41% 5.35
+SimPO 49.31 5.2% 5.8% 6.0% 3.3 119.84 9.59% 13.58% 15.4% 5.63
+TDPO 83.1 6.2% 6.58% 6.8% 3.26 204.6 9.23% 10.86% 10.7% 5.55
+SparsePO 49.3 4.62% 6.17% 6.35% 3.0 85.0 9.76% 13.58% 12.8% 5.43
+SePO-rand 24.92 3.07% 4.26% 4.4% 2.86 68.3 6.73% 6.38% 6.47% 5.23
+SePO (Ours) 29.84 6.3% 7.1% 7.32% 3.45 71.97 10.3% 14.4% 14.91% 6.38
Base Pythia-6.9B 4.23% 5.0% 5.17% 3.58 LLaMA2-Chat-13B 12.0% 8.4% 7.7% 5.7
+DPO 190.9 10.2% 12.78% 13.27% 4.7 319.67 13.48% 13.72% 13.37% 5.84
+IPO 184.74 8.1% 11.78% 12.6% 4.34 336.33 13.95% 14.27% 14.4% 5.76
+RRHF 141.27 7.47% 11.42% 13.2% 4.31 269.61 13.84% 15.94% 16.36% 5.73
+SimPO 145.26 8.0% 11.8% 12.72% 4.51 270.35 14.7% 16.4% 17.02% 5.7
+TDPO 209.93 10.68% 13.92% 13.7% 4.78 297.43 14.4% 15.0% 15.65% 6.37
+SparsePO 96.43 8.6% 12.45% 11.9% 4.52 161.0 13.7% 15.6% 15.32% 5.84
+SePO-rand 77.35 4.82% 5.28% 5.46% 3.45 140.57 10.05% 8.16% 7.5% 5.66
+SePO (Ours) 79.07 10.94% 14.27% 13.6% 5.09 151.05 15.5% 17.53% 18.41% 6.86
Base LLaMA3-Base-8B 3.3% 6.2% 4.6% 5.4 LLaMA3-Instruct-8B 20.60% 22.9% 22.6% 6.5
+DPO 210.0 15.9% 18.2% 15.5% 6.36 201.0 32.6% 40.3% 37.9% 6.94
+IPO 215.73 17.8% 14.4% 14.2% 6.7 212.46 30.5% 35.6% 35.6% 6.94
+RRHF 148.8 6.3% 12.1% 10.1% 5.8 162.44 26.5% 34.4% 36.0% 6.7
+SimPO 128.1 23.4% 22.0% 20.3% 6.76 136.9 33.8% 50.0% 48.8% 7.03
+TDPO 245.8 24.7% 23.9% 22.0% 5.94 245.52 29.7% 45.0% 37.6% 6.53
+SparsePO 94.0 20.6% 21.78% 21.02% 6.4 96.4 31.1% 44.34% 41.0% 6.7
+SePO-rand 83.9 6.2% 10.6% 8.3% 5.6 89.23 25.5% 26.74% 25.5% 6.2
+SePO (Ours) 84.5 25.0% 25.6% 22.78% 6.8 84.36 34.35% 50.5% 47.04% 6.94
Table 1: Main results. For SePO, the oracle models are fully trained TinyLLaMA-1.1B and Pythia-1B, and used to select the top-30% tokens. "GPU Hours" records the GPU running time of each method on different policy models.
@@ 4.1 Experimental Settings
Models and Training Data.
To approximate the optimal token-level reward function, we first obtain the reference models by training on UltraChat-200K (Ding et al., 2023) in an SFT manner. We train reference models on the TinyLLaMA-1.1B (Zhang et al., 2024) and Pythia-(70M, 160M, 410M, 1B, 1.4B) (Biderman et al., 2023) to facilitate research on the effect of oracle models with different sizes. For each reference model, we obtain the oracle models by further fine-tuning with DPO on UltraFeedback (Cui et al., 2023). To obtain target policy models, we perform SFT on 3 foundation models: Pythia-(2.8B, 6.9B) and LLaMA3-Base-8B (Dubey et al., 2024), with UltraChat-200K. We also test SePO on well-aligned models LLaMA2-Chat-(7B,13B) (Touvron et al., 2023), and LLaMA3-Instruct-8B.
Baseline Methods.
We compare SePO with state-of-the-art offline preference optimization methods: DPO (Rafailov et al., 2024c), IPO (Azar et al., 2024), RRHF (Yuan et al., 2024) and SimPO (Meng et al., 2024), and token-level alignment method TDPO (Zeng et al., 2024) and SparsePO (Christopoulou et al., 2024) (learnable sparse mask). To evaluate the SePO token selection, we further include a self-implemented SePO-rand baseline that randomly selects k% tokens from the pair-wise data and optimizes via Eqn. 11.
Evaluation Benchmarks.
We quantify the results by evaluating on three widely used instruction-following benchmarks: AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2024b), and Arena-Hard (Li et al., 2024). All judgments are performed by the latest GPT-4o model. More details in Appendix G.
@@ 4.2 Overall Performance
Performances of SePO and other baseline methods on three benchmark datasets are presented in Table 1. According to the results, SePO significantly improves performance over the base policy models, with an average of 9.05% improvement in win rates on Arena-Hard. SePO also outperforms other strong preference optimization methods. On MT-Bench, SePO achieves the best average scores among other methods on five of six policy models, surpassing both state-of-the-art response-level methods such as SimPO and token-level method TDPO and SparsePO. Notably, SePO models are only optimized on 30% of the tokens trained on other methods, leading to 39.94%-62.34% reduction in GPU training hours compared to baseline methods. Further discussions on how SePO saves GPU training hours and the influence of oracle model training on SePO efficiency are shown in Appendix E. These results directly strengthen the effectiveness of selective training strategies applied in preference optimization. Further comparisons with SePO-rand show that optimizing on randomly selected k% tokens significantly damages the performance of selective training, proving the effectiveness of our DPO-based token selection strategy in filtering the crucial supervision signals from the training data.
On AlpacaEval 2.0, SePO continues to achieve superior performance over baseline methods in both win rates and LC win rates. Notably, SePO outperforms all other methods on length-controlled win rates, including SimPO and RRHF, which are specifically designed for length-normalized reward formulation. These results show that selective training strategies also enhance policy models in avoiding over-length responses. We believe the reason is that during token selection, the token-level reward function can assign the end-of-sentence tokens with low-ranking scores, which can be discarded during optimization if the response ends inappropriately (e.g. over-length or repetition). In contrast, though SimPO and RRHF design length-normalized rewards, the end-of-sentence tokens are still included and fitted for all training samples.
@@ 4.3 Impact of Data Scale
Figure 3: SePO with different combinations of K% selection ratios for chosen/rejected responses, quantified by the LC win rates on AlpacaEval 2.0.
According to the results, increasing selection rates from 0.1 for chosen/rejected responses rapidly improves policy model performance, but the momentum decreases as the ratio continues to grow. For example, the LC win rate of LLaMA2-Chat-7B improves from 8.37% to 14.8% as the ratios for chosen/rejected responses rise from 0.1 to 0.5 progressively, then stabilizes around 14.7% with higher selection rates. These observations prove our hypothesis that not all tokens are equally effective for LLM alignment. Training only on key tokens effectively improves alignment performance, while other tokens provide limited supervision information. Training on Top-50% tokens for LLaMA2-Chat-(7B, 13B) provides comparable performance to aligning on all tokens.
Figure 4: LC win rates on AlpacaEval 2.0, supervised by oracle models trained with different data proportions. We report the average performance of 3 random runs.
Figure 5: (a) LC win rates on AlpacaEval 2.0, trained with oracle models of various sizes; (b)(c) token-level reward distributions for 5,000 chosen/rejected responses obtained from oracle models with different sizes.
@@
In Theorem 2, we proved that training an oracle model on a moderate-scale subset is a pessimistic estimation of the target reward function. In this section, we empirically investigate the influence of the training data scale ($\frac{m}{N}$ in Theorem 2) for oracle models. Specifically, we randomly sample different proportions of data from UltraFeedback as training data for the TinyLLaMA-based oracle model. The results are shown in Figure 4.
@@ 4.4 Weak-to-Strong Generalization
In Table 1, we presented the performance of weak oracle models on guiding stronger policy models (e.g. LLaMA2-Chat-13B). The competitive results of SePO prove the viability of weak oracle modeling. To provide a clear landscape, we further train oracle models with Pythia-(70M, 160M, 410M, 1B, 1.4B) on the same target data and compare their performances on the Pythia-SFT-(2.8B, 6.9B) policy models. The results are shown in Figure 5 (a).
[…]
To provide an intuitive view, we present the token-level score distributions of different oracle models in Figure 5 (b)(c). For both chosen/rejected scores distribution, strong oracle models such as Pythia-(1B,1.4B) show higher densities in extreme (large and small) reward values, which facilitates separating key tokens from the other tokens. In contrast, small oracle models tend to fit a Gaussian distribution, where most tokens have similar scores. These results show that strong oracle models excel in distinguishing key tokens within the dataset, which further proves the capability of the oracle models as crucial in accurately modeling the token-level reward function.
Weak Data Supervision.
We evaluate the weak data supervision performance of SePO by training on HH-RLHF (Bai et al., 2022), an early-released preference dataset with relatively lower quality (Yang et al., 2024a) on responses. We perform SePO with a TinyLLaMA-based oracle model and 30% token selection rates, and comparisons with baseline methods are shown in Table 2. According to the results, SePO is the only method that improves the strong LLaMA2-Chat-13B policy model with data from HH-RLHF, outperforming base performance by 1.63% on Arena-Hard and 0.41% on AlpacaEval 2.0. With full optimization, baseline methods such as DPO and SimPO continuously degrade model performance due to over-optimization on weak supervision data. These results prove SePO effective in leveraging useful supervision signals from weak data while avoiding over-fitting harmful patterns. These results point to SePO as a highly efficient method for continually improving strong model performance with large-scale out-of-distribution data.
@@
Table 2: Weak data supervision performance.
4.5 Scaling Supervision for SePO
To evaluate the boundary of SePO capability, we further scale the size of the oracle model to provide strong supervision beyond target model capabilities. We refer to the results in Figure 5 (a) for experimental results from the oracle model size 70M-1.4B, measured in performance on AlpacaEval 2.0. The performance shows a scaling trend as the oracle model size increases. In Table 3, we further extend the oracle model size to 12B on supervising Pythia-(2.8B,6.9B) policy models on the same target data.
Oracle Model Size Pythia-2.8B Pythia-6.9B
1B 7.1% 14.27%
1.4B 8.07% 15.21%
2.8B 9.4% 16.24%
6.9B 9.85% 16.8%
12B 10.1% 17.14%
Table 3: LC win rates on AlpacaEval 2.0, trained with scaling oracle model sizes from 1B to 12B.
According to the results, the policy models’ performance continually improves as the oracle model size increases, with about 3% improvement on both policy models from a 1B oracle model to a 12B oracle model. These results show that the performance of SePO can further improve along with oracle models that better estimate the token selection function. However, we also observe an overall trend in decreasing momentum of the improvement as the oracle model size increases, indicating a potential information bottleneck for the selective training paradigm. These results show that the use of stronger oracle models can significantly increase SePO costs while providing a limited performance improvement.
4.6 Related Work
Previous works related to SePO can be divided into three parts: response-level preference optimization (Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2024c; Yuan et al., 2023; Meng et al., 2024; Ethayarajh et al., 2024; Lu et al., 2024; Azar et al., 2024), token-level optimization (Rafailov et al., 2024b; Zhong et al., 2024; Zeng et al., 2024; Yoon et al., 2024; Chen et al., 2024c; Chan et al., 2024; Lai et al., 2024; Chen et al., 2024b), and weak-to-strong generalization (Burns et al., 2023; Lang et al., 2024; Yang et al., 2024b; Charikar et al., 2024; Zhou et al., 2024b; Ji et al., 2024; Zheng et al., 2024a). We provide a detailed description of related work in Appendix A.
5 Conclusion